### v1

# Modular Jump Gaussian Processes

Abstract
Gaussian processes (GPs) furnish accurate nonlinear predictions with well-calibrated uncertainty. However, the typical GP setup has a built-in stationarity assumption, making it ill-suited for modeling data from processes with sudden changes, or “jumps” in the output variable. The “jump GP” (JGP) was developed for modeling data from such processes, combining local GPs and latent “level” variables under a joint inferential framework. But joint modeling can be fraught with difficulty. We aim to simplify by suggesting a more modular setup, eschewing joint inference but retaining the main JGP themes: (a) learning optimal neighborhood sizes that locally respect manifolds of discontinuity; and (b) a new cluster-based (latent) feature to capture regions of distinct output levels on both sides of the manifold. We show that each of (a) and (b) separately leads to dramatic improvements when modeling processes with jumps. In tandem (but without requiring joint inference) that benefit is compounded, as illustrated on real and synthetic benchmark examples from the recent literature.
Keywords: computer experiment, nonstationarity, latent variable, local neighborhoods, mixture model
1 Introduction
We consider processes $Y(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ that are smooth for most of the surface except along a manifold of discontinuity which spatially partitions the response $Y(x)$ into two levels depending on $x$. One example of such a process is a computer simulation measuring the average transport time $Y(x)$ of material transport devices given their characteristics $x$ (Park et al.,, 2023). Usually, transport time is low, but when the devices are working at or near capacity transport time suddenly increases – it “jumps”. Other examples of jump dynamics exist in geostatistics (Kim et al.,, 2005), physiology (Vieth,, 1989), ecology (Toms and Lesperance,, 2003), and materials science (Park et al.,, 2022). Our interest is in jumps exhibited by computer simulation experiments, where $Y(x)$ models complex phenomena as a deterministic function of inputs $x$ at potentially great computational expense. In this paper we consider jump processes with only two levels.
When designing computer simulation experiments, one often meta-models a planned experiment of runs of the computer model. That is, gather a collection of $N$ data pairs $(x_{1},y_{1}),\dots,(x_{N},y_{N})$, where $y_{i}=Y(x_{i})$, and obtain a fit that can predict $Y(x)$ at new inputs. Such surrogate models (Santner et al.,, 2003; Gramacy,, 2020) are instrumental to many applications where ready access to runs of the computer model are unavailable or expensive. The canonical surrogate is a Gaussian process (GP; e.g., Sacks et al.,, 1989) because GPs furnish accurate nonlinear predictions with good uncertainty quantification. For this reason they are also popular for many regression applications in machine learning (e.g., Rasmussen and Williams,, 2006) and in geospatial settings (e.g., Banerjee et al.,, 2004). However there are two downsides to GP modeling in many contexts. One is that computational demands for inference scales poorly as $N$ gets large; another is that the typical GP formulation assumes stationarity, in that only relative distances between input locations matter. This means GPs struggle to capture dynamics where modeling fidelity varies in the input space, e.g., where one part is flat and another is wiggly. GPs also struggle if there are jumps of discontinuity in which the level of the response drastically changes.
There is a sizable literature on nonstationary GP modeling. Sauer et al., 2023a provide a recent review. Many such approaches involve imposing hard partitions in the input space (e.g., Kim et al.,, 2005; Pope et al.,, 2021; Payne et al.,, 2020; Gramacy and Lee,, 2008; Luo et al.,, 2023), which could potentially cope with jumps that are axis aligned (like with trees) or are polygonal (Voronoi tesselations). Other approaches involve smoothly warping the inputs into a stationary regime (e.g., Sampson and Guttorp,, 1992; Schmidt and O’Hagan,, 2003; Damianou and Lawrence,, 2013; Sauer et al., 2023b, ), which is often better but would struggle with discontinuity. GPs exist which can handle discontinuities (e.g., Garnett et al.,, 2010; Picheny et al.,, 2019), but they do not assume the response contains distinct levels. Park, (2022) created the “jump GP” (JGP) specifically for data from processes that are piecewise continuous in order to address this gap. JGPs have three main modeling ingredients that are tightly coupled together. The first two are local (GP) modeling (e.g., Gramacy and Apley,, 2015), and latent component (i.e., output level) membership variables. Inference is joint over all unknowns – in particular for the high dimensional latent variable – via modern variants of expectation maximization (EM; Dempster et al.,, 1977), comprising the third ingredient. JGP is a clever, but complex, modeling apparatus that works well out-of-sample when trained on data from response surfaces exhibiting jumps.
Here we suggest that the essence of a JGP can be extracted by appropriating its main ingredients in a more modular setup. We still use local modeling, latent variables, and clustering via EM, but we avoid joint inference and instead do things one step at a time by daisy-chaining off-the-shelf subroutines. To explain, consider a simple illustrative example of a process with jumps in 1d.
$Y(x)=\begin{cases}\sin(x)&x<50\\ \sin(x)+10&50\leq x<70\\ \sin(x)&x\geq 70\end{cases}$ (1)
So $Y(x)$ follows a simple sinusoid but jumps up by 10 between $x=50$ and $x=70$. Each panel of Figure 1 shows $Y(x)$ as a solid black line with a different “fit” from data $(x_{1},y_{1}),\dots,(x_{N},y_{N})$. We use $N=100$ inputs $x_{i}$ chosen via Latin hypercube sample (LHS; McKay et al.,, 1979). These are not shown to reduce clutter.
Refer to captionRefer to captionRefer to captionRefer to caption +Local +Feature +Local +Feature
Figure 1: All panels show true function $Y(x)$ from Eq. (1) in black. Colored lines represent the mean (solid) and 95% PI (dashed) via (top left) a Global GP, (top right) LAGP, (bottom left) Modular JGP with global GP, and (bottom right) Modular JGP with LAGP. Training data not shown.
Figure 1 was produced with standard libraries and default settings, detailed in Section 5.
The top left panel of Figure 1 shows an ordinary/stationary GP fit, as reviewed in Section 2.1. Three issues are of note here: the GP is mistaking signal for noise, overestimating variance, and smoothing over the jumps. It seems to only have detected signal from the jumps, not from the sinusoid. Uncertainty is uniform because inputs are uniform and the GP assumes stationarity. Now turn to the top right panel, showing a local (approximate) GP (LAGP), reviewed in Section 2.2. For now we remark that a LAGP is basically an ordinary GP where training data far from the testing location(s) is not used. Less data means faster calculations and increased “reactivity” if dynamics change in the input space. Notice how LAGP shows an improvement over the ordinary GP (top left) for $x<40$ and $x>80$. Signal is detected there, and variance is much reduced. But this comes at the expense of accuracy in the middle region, and near the “jumps” on either side. Uncertainty here is inflated relative to the stationary GP fit. Importantly, LAGP uses a fixed-sized neighborhood of locality everywhere in the input space which we shall explain in more detail in Section 2.2. Our proposed approach improves upon this by dynamically selecting the optimal neighborhood size throughout the input space as described in Section 3.
The lower panels of Figure 1 highlight our main contribution. Focus first on the lower-left. Suppose we were to cluster the observations marginally. We ignore $(x_{1},\dots,x_{N})$ and use an unsupervised learning scheme to assign (one of two) labels to each of $(y_{1},\dots,y_{N})$. In this 1d example, most clustering methods will easily partition these $y$-values into components around 0 and 10. We use model-based clustering via EM for this, as described in Section 4.2, but suggest several variations and other approaches that often work just as well. Cluster memberships in hand, e.g., $c_{i}=1$ for $y_{i}$ around 0 and $c_{j}=2$ for $y_{j}$ around 10, suppose we were to learn a classification rule spatially for those pairs $(x_{1},c_{1}),\dots,(x_{N},c_{N})$, i.e., bringing the $x$’s back in with supervised learning. There are again many choices for this; we prefer a classification GP (CGP). This allows us to predict a probability of class 2 at $x$, in-sample at ($x_{1},\dots,x_{N}$) yielding $(p_{1},\dots,p_{N})$, or out-of sample for testing as $p(x)$. We then treat the class predictions $(p_{1}\dots p_{N})$, derived from the training data, as an input feature in our GP fit. That is, take the same GP from the top-left panel, but train it on input tuple $(x_{i},p_{i})$ and output $y_{i}$, for $i=1,\dots N$. When predicting, augment any testing $x$ with $p(x)$ for tuple $(x,p(x))$. This global GP with the new feature is visualized in the lower left panel of Figure 1. Notice how the fit tracks the jumps better (compared to the top left), and it is more confident everywhere, but it still struggles to pick up on the sinusoid. The idea of learning an auxiliary input feature to accommodate nonstationarity is not new (Bornn et al.,, 2012) but, like Park et al., (2022) for JGP, Bornn et al., perform joint inference for the GP and the latent quantity, which can be cumbersome. A key feature of our approach is that we do everything in a pre-processing step with off-the-shelf libraries. This both increases the flexibility of our approach and makes it more widely accessible.
Finally, the bottom right panel of Figure 1 combines the feature with local modeling: LAGP applied on $(x_{i},p_{i})$ and output $y_{i}$, but otherwise the same as the feature-expanded GP on the left. We call this the Modular JGP. Except for some elevated uncertainty near the jump points, which is sensible, the fit is nearly perfect: high accuracy, low variance. In the remainder of the paper our goal is to flesh out these ideas, making them more concrete, especially as regards implementation. We begin with review in Section 2. We then describe our main methodological contributions in Sections 3 and 4: optimizing local neighborhoods and cluster-based features. Throughout, we provide illustrations on examples that have the same spirit as Figure 1, but are more realistic. Section 5 provides empirical benchmarking exercises in several variations including on a real simulation from manufacturing. Finally, Section 6 concludes with a brief discussion of limitations and possible future work. All code is provided via Git: https://bitbucket.org/gramacylab/jumpgp.
2 Review
Here we provide a brief review of GPs and a popular variant, LAGP.
2.1 Gaussian Processes
Let $X_{N}$ be an $N\times d$ matrix of input features and $Y_{N}$ be a corresponding $N\times 1$ vector of outputs. Additionally, let $x_{i}^{\top}$ denote the $i^{\mathrm{th}}$ row of $X_{N}$. We consider estimating an unknown function via training data $X_{N}$ and $Y_{N}$ using a GP (Rasmussen and Williams,, 2006; Gramacy,, 2020). A typical GP prior implies that $Y_{N}\sim\mathcal{N}_{N}\bigl{(}\mu,\Sigma(X_{N})\bigr{)}$, where $\mu$ is often assumed to be 0 so that all of the modeling “action” is in $\Sigma(X_{N})$. This setup induces a likelihood function:
$L(Y_{N}\mid X_{N})\propto\bigl{|}\Sigma(X_{N})\bigr{|}^{-\frac{1}{2}}\mathrm{exp}\left\{-\frac{1}{2}Y_{N}^{\top}\Sigma(X_{N})^{-1}Y_{N}\right\}.$ (2)
The likelihood (2) can be used to learn $\Sigma(X_{N})$ under some restrictions, such as positive definiteness. Often, covariance built from a kernel specifying $\Sigma(X_{N})_{ij}=\Sigma(x_{i},x_{j})$. Kernels that depend only on the distance between inputs yield stationary processes, eliminating all input position information from the model. One example is squared exponential kernel
$\Sigma(X_{N})_{ij}=\Sigma(x_{i},x_{j})=\tau^{2}\mathrm{exp}\left\{-\sum_{k=1}^{d}\frac{\left\|x_{ik}-x_{jk}\right\|^{2}}{\theta_{k}}\right\},$ (3)
which we use throughout, although none of our contribution is explicitly tied to this choice. Other stationary kernels, such as Matérn (Stein,, 2012), are similarly appropriate. Above, $\tau^{2}$ and $\theta=[\theta_{1},\dots,\theta_{d}]$ are hyperparameters representing the scale and lengthscale, respectively. These quantities may be inferred via maximum likelihood estimation (MLE) with Eq. (2).
The same structure may be extended from training to testing for prediction. Let $\mathcal{X}$ denote an $m\times d$ matrix of new/testing locations, and $\tilde{x}^{\top}_{i}$ be the $i^{\text{th}}$ row of $\mathcal{X}$. Then, using any hyperparameters learned from training data, $Y(\mathcal{X})\mid X_{N},Y_{N}\sim\mathcal{N}_{m}\left(\mu_{N}({\mathcal{X}}),\Sigma_{N}({\mathcal{X}})\right)$ where
$\mu_{N}({\mathcal{X}})$ $=\Sigma(\mathcal{X},X_{N})\Sigma(X_{N})^{-1}Y_{N},$ using $\mbox{using }\;\Sigma(\mathcal{X},X_{N})_{ij}$ $=\Sigma(\tilde{x}_{i},x_{j}),$
and $\mbox{and }\quad\Sigma_{N}({\mathcal{X}})$ $=\Sigma(\mathcal{X})-\Sigma(\mathcal{X},X_{N})\Sigma(X_{N})^{-1}\Sigma(\mathcal{X},X_{N})^{\top}.$ (4)
As a shorthand for later, let $\mathrm{GP}\left(\mathcal{X};X_{N},Y_{N}\right)$ indicate the use of Eqs. (24) to represent predictions at $\mathcal{X}$ via MLE hyperparameters using training data $X_{N}$ and $Y_{N}$. We refer to this as a “global GP” since all training data are used. Such a fit is valid for prediction at any $\mathcal{X}$ in the input space. This is in contrast to a GP fit using a subset of the training data, which we shall review next. The GP shown in the top left panel of Figure 1 is a global GP, where the solid yellow line is $\mu_{N}(\mathcal{X})$ and the dashed yellow line is a 95% prediction interval (PI) calculated using the diagonal entries of $\Sigma_{N}(\mathcal{X})$.
A global GP has two drawbacks: computation becomes prohibitive as $N$ increases, and it does not perform well when training data come from a nonstationary processes. The first can be seen by noting $\Sigma(X_{N})^{-1}$ and $|\Sigma(X_{N})|$ in Eqs. (2) and (4), both $\mathcal{O}(N^{3})$ operations. To see the second, refer back to the discussion of Figure 1 (top-left), illustrating fits on a process with jumps. Consider a single data point at $x=45$, which is part of the level-0 region. The issue with the global GP setup is that two equidistant points, say $x=35$ and $55$, generate the same covariance, even though one point is in the level-0 region and the other is in the level-10 region. However notice that the data is locally stationary, in three spatial regimes, which suggests a divide-and-conquer approach may work well.
2.2 Local Approximate Gaussian Processes
LAGP (Gramacy and Apley,, 2015) tackles both drawbacks at once: computational expense and stationarity. Observe from Eq. (3) that for two locations $x_{i}$ and $x_{j}$, $\Sigma(x_{i},x_{j})\rightarrow 0$ as $\left\|x_{i}-x_{j}\right\|\rightarrow\infty$. Consider a single new location $x$ where we wish to make a prediction, e.g., a single row of $\mathcal{X}$ in Section 2.1. Note in Eq. (4) that $\Sigma(x,X_{N})$ serves as a weight for $\mu_{N}(\mathcal{X})$; thus, any training $y_{i}$ whose $x_{i}$ is far from $x$ has little influence on $\mu_{N}(\mathcal{X})$. Recognizing this, LAGP chooses a subset of $n\ll N$ “useful” data points $X_{n}(x)\subset X_{N}$ to predict $x$. There are many ways to define “useful”, such as nearest neighbors (NN), mean-squared predictive error, and Fisher information. Throughout we consider LAGP via NN, although it could be worth entertaining other criteria as we discuss briefly in Section 6.
The core idea behind LAGP is that fitting a GP to $X_{n}(x)\subset X_{N}$ with $n\ll N$, along with corresponding $Y_{n}$, is both faster, and more reactive to local dynamics around $x$. We encapsulate that in shorthand as $\text{LAGP}_{n}(x)=\text{GP}\left(x;X_{n}(x),Y_{n}\right)$ after optimizing any hyparameters in $\Sigma(X_{n}(x))$, and $\text{LAGP}_{n}(\mathcal{X})$ for $m$ separate applications covering all of $\mathcal{X}$. The top right panel of Figure 1 shows $\text{LAGP}_{25}(\mathcal{X})$, where the red solid line denotes $\mu_{N}(\mathcal{X})$ and the red dashed line denotes a 95% PI. Since calculations for each ${x}\in\mathcal{X}$ are statistically independent, we don’t get a full covariance structure $\Sigma_{N}(\mathcal{X})$, only a diagonal one with entries $\sigma^{2}_{x}$. The upside, though, is that the calculations are massively parallelizable (Gramacy et al.,, 2014).
LAGP reduces computational costs from $\mathcal{O}(N^{3})$ to $\mathcal{O}(n^{3})$, offering substantial speedups if $n\ll N$. Much work has been done to further decrease the computation time for fixed neighborhood size $n$ (Gramacy and Haaland,, 2016; Sung et al.,, 2018; Cole et al.,, 2021). Besides being faster, LAGP may also be more accurate for data from a nonstationary process. To see this, refer again to Figure 1 (top-right). By excluding points that are far from each $x$, LAGP has captured local trends (at least for the $x<40$ or $x>80$ regions) better than the global GP (top-left). However, LAGP does not perform well in the $8040<x<$ 80 region. This is because these predictive locations have neighbors from across the jumps.
To explore a potential remedy, consider a higher-dimensional illustration.
Figure 2: Three possible neighborhood choices for $x=(0.04,0.023)$ (the blue dot). The two colors (red and yellow) are two levels of the response, with a manifold of discontinuity at $\sin(x_{1})=x_{2}$. The orange and yellow dots are $X_{N}$ and the black dots are three different choices for $X_{n}(x)$.
Figure 2 depicts a simple manifold of discontinuity with two output values, orange and yellow, observed on an $X_{N}$ grid. Now consider predicting at $x=(0.04,0.23)$, in blue, using LAGP. Some potential neighborhoods $X_{n}(x)$ are depicted in black, one in each panel. The left panel uses Euclidean NNs with $n=50$, but many such neighbors come from the opposite side of the manifold. One workaround is to shrink the neighborhood size, e.g., $n=13$ in the center panel, which is the largest $n$ that only contains neighbors from the orange region. So the choice of $n$ implies a tradeoff between accuracy and reactivity. It could be advantageous to have a larger neighborhood (e.g., $n=50$), but comprised of only orange points, as shown in the right panel.
3 Optimal Neighborhood-size LAGP
We turn now to the neighborhoods depicted in the center [this section] and right [Section 4] panels of Figure 2, respectively. The idea is to make adjustments near a manifold of discontinuity without affecting modeling dynamics elsewhere. To aid illustrations, Figure 3 borrows an example from Park, (2022), known as “Phantom” data, with details in Appendix A.1.
Figure 3: Left: True “Phantom” with training grid in gray; Right: Histogram of the marginal response.
The image on the left is formed by evaluating the test function on a dense $101\times 101$ grid in $[0,1]^{2}$. All responses ($y$-values) are deterministic. Observe that there is smooth color variation within the (mostly) red and yellow regions, revealing a strong signal locally (light reds and dark yellows, respectively), even away from the “jumps”. The histogram on the right helps understand the magnitudes of these signals. Notice that red and yellow regimes are have comparable within-color variability and they are well-separated, with no values between 0.4 and 0.6.
The visuals in the panels of Figure 2 could represent zoomed-in versions of any of the jumps separating red from yellow in the left panel of Figure 3. Here, in Section 3, we present the first prong, which targets finding the optimal $n(x)$ for LAGP [Figure 2, middle] using its default, radial topography. The right panel of Figure 2, morphing that topography, represents the second prong and is the subject of Section 4. While jumps naturally focus interest on inputs near the manifold, the value of dynamically determining $n(x)$ for LAGP (other than the software default of $n=50$) is not limited to those areas only. Intuitively, regions of slowly varying response should benefit from larger neighborhoods, whereas ones where it is changing more rapidly (via jump or not) should benefit from the greater reactivity provided by a smaller one.
3.1 Searching via validation
Consider a single predictive location $x$ and a set of possible neighborhood sizes $n\equiv n(x)\in\{n_{\min},\dots,n_{\max}\}$. We wish to estimate an optimal $\hat{n}(x)$ by some criterion, where $X_{n_{\min}}(x)\subseteq X_{\hat{n}}(x)\subseteq X_{n_{\max}}(x)\subseteq X_{N}$, as solved by LAGP [Section 2.2]. Let $\mu_{n}(x)$ denote the predictive mean under some $X_{n}(x)$, i.e., following Eq. (4), but conditional on $(X_{n}(x),Y_{n})$. This prediction has some accuracy, say measured by squared error $\mathrm{SE}(y,\mu_{n}(x))$ where $y$ is the true, but unknown value of the response at $x$. Generically, $\mathrm{SE}(y,\hat{y})=(y-\hat{y})^{2}$. One of the main contributions of Gramacy and Apley, (2015) involves arguing that NN-constructed $X_{n}(x)$, conditional on a neighborhood size $n$, approximately minimizes the the expectation of this quantity (i.e., the mean-squared error) for $y$ following a stationary GP.
Here, we wish to extend that to choosing the best $\hat{n}$. However, we can’t use MSE under a stationary GP for $y$. There are two reasons for this. One is practical: this is not the case we’re interested in – our jump processes are nonstationary. The other is more fundamental. If the process is stationary, then MSE is a decreasing function of $n$: you want as much training data as possible. Since large $n$ is generally intractable, Gramacy and Apley, recommended choosing $n$ as large as computational constraints allow. However, they observed that performance varied with $n$ when the data exhibited telltale signs of nonstationarity, leaving an open question about what to do in practice. So instead of averaging SE, we create a validation exercise which allows us to plug-in an actual $y_{i}$ value from the training data as a proxy.
Let $x_{*}$ be the input in $X_{N}$ that is closest (in terms of Euclidean distance) to $x$. Note that this is also the closest point to $x$ in any $X_{n}(x)$ for $n\geq 1$ and that $x_{*}\equiv X_{1}(x)$. We prefer the simpler $x_{*}$, which is easier on the eyes than $X_{1}(x)$, and allows us to use $y_{*}$ to notate the corresponding element of $Y_{N}$, our proxy for the unknown $y(x)$. Next, let $X_{n}(x_{*})$ denote an $n$-sized neighborhood for predicting at $x_{*}$, i.e., the $n$-nearest points to $x_{*}$ in $X_{N}$, excluding $x_{*}$. Using these quantities, and for any neighborhood size $n$, one may fit $\text{LAGP}_{n}\left(x_{*};X_{n}(x_{*}),Y_{n}\right)$, furnishing a prediction for $y_{*}$ via $\mu_{n}(x_{*})$, streamlining notation. Suppose we do this for each $n$ of interest, and take the best one by squared error:
$\hat{n}(x_{*})=\underset{n\in\{n_{\min},\dots,n_{\max}\}}{\arg\min}\mathrm{SE}\left(y_{*},\mu_{n}(x_{*})\right).$ (5)
This $\hat{n}(x_{*})$ is the best neighborhood size for $x_{*}$, which is nearby the location of interest $x$. What remains is to refit the local model at $x$ using the $\hat{n}(x_{*})$-nearest neighbors to $x$, $\mathrm{LAGP}_{\hat{n}(x_{*})}(x;X_{\hat{n}}(x),Y_{\hat{n}})$.
Without care, fitting $(n_{\max}-n_{\min})$-many GPs in pursuit of Eq. (5) could be prohibitively expensive. One option is parallelization, since the calculation for each $n$ is statistically independent, and thus computationally exclusive, of the others. However, we prefer to parallelize instead over the predictive location(s), $x\in\mathcal{X}$, which will be described in Section 5. A computationally thrifty solution takes advantage of the nested structure of neighborhoods as $n$ increases: $X_{n_{\min}}(x_{*})\subseteq X_{n_{\min}+1}(x_{*})\subseteq\dots\subseteq X_{n_{\max}}(x_{*})$. Observe that any two neighborhoods $X_{n}(x_{*})$ and $X_{n-1}(x_{*})$ are identical except for one point: the $n^{\mathrm{th}}$ furthest point from $x_{*}$. Let $x_{n}=X_{n}(x_{*})\setminus X_{n-1}(x_{*})$ denote this new location as the neighborhood increases by one. A visual of these $x_{n}$, selected from $X_{N}$ on a $25\times 25$ grid, is provided in the left panel of Figure 4.
Figure 4: Left: Order of LAGP neighborhood for point $(-0.32,0.17)$; Center: SE by neighborhood size for point $(-0.32,0.17)$; Right: Optimal neighbors LAGP (OLAGP) fit. Blue dots indicate two predictive sites and the surrounding black dots are the respective optimal neighborhoods.
Each number indicates the position of $x_{n}$ for some $n$ relative to $x_{*}\in X_{N}$, the closest point to $x=(-0.32,0.17)$.
Now, proceeding in order for $n=n_{\min},\dots,n_{\max}$, suppose we had already calculated $\mathrm{LAGP}_{n-1}(x_{*})$, beginning initially with $n-1\equiv n_{\min}$ via the description in Section 2.2, yielding $\Sigma_{n-1}^{-1}\equiv\Sigma(X_{n-1}(x_{*}))^{-1}$ and similarly for $\log|\Sigma_{n-1}|$. Incorporating the next point, $x_{n}$, involves adding a new row/column to $\Sigma_{n-1}$ to get $\Sigma_{n}$, which may be calculated without new, cubic cost, matrix decompositions by following the partition inverse equations (Barnett,, 1979). Specifically,
$\Sigma_{n}^{-1}$ $=\begin{bmatrix}[\Sigma_{n-1}^{-1}+hh^{\top}v]&h\\ h^{\top}&v^{-1}\end{bmatrix}$ and $\log|\Sigma_{n}|$ $=\log|\Sigma_{n-1}|+\log(v),\quad\text{where}$ (6)
$h$ $=-v^{-1}\Sigma_{n-1}^{-1}\Sigma(X_{n}^{-1}(x_{*}),x_{n}),$ and $v$ $=\Sigma(x_{n})-\Sigma(x_{n},X_{n-1}(x_{*}))\Sigma_{n-1}^{-1}\Sigma(X_{n-1}(x_{*}),x_{n}).$
Each update is quadratic in $n$, so that the entire processes require $c\sum_{n}^{n_{\max}}n^{2}\in\mathcal{O}(n_{\max}^{3})$ flops. In other words, the computational expense is no greater than that involved in computing the largest LAGP in the search. Given these quantities, $\mu_{n}(x_{*})$ required as in Eq. (4) is again at worst cubic in $n$.
These updates, and the computational cost analysis, presume that $\Sigma_{n-1}$ and $\Sigma_{n}$ utilize the the same hyperparameters, like lengthscales $\theta_{k}$, which may not be ideal. In particular, shorter lengthscales will generally be preferred for smaller $n_{\min}$-sized neighborhoods whose squared distances (3) are short. Whereas when $n$ is closer to $n_{\max}$, including longer squared distances, longer lengthscales may be preferred. However, an MLE calculation (2) after each update, as new $x_{n}$ arrives, would incur cubic costs and result in a quartic overall computational complexity, which isn’t ideal either. Recognizing that an additional data point is of greater value to the likelihood, say via its Fisher information, when $n$ is small as opposed to large, we deploy a logarithmic MLE-updating schedule so that cubic MLE costs are incurred much less often as $n$ increases. Any sub-linear updating schedule will keep computation cubic.
Algorithm 1 OLAGP (Optimal Neighbors LAGP)
1:$X_{N},Y_{N},x,n_{\min},n_{\max}$
2:$x_{*}=X_{1}(x)$triangleright nearest neighbor of $x$
3:$\{\Sigma_{n_{\min}}^{-1},\log|\Sigma_{n_{\min}}|\}=\mathrm{chol}(\Sigma(X_{n_{\min}}(x_{*})))$triangleright calculate $\Sigma_{n_{\min}}^{-1}$ and $\log|\Sigma_{n_{\min}}|$ using (3)
4:$\{\tau^{2},\theta\}=\mathrm{MLE}(\Sigma_{n_{\min}}^{-1},\log|\Sigma_{n_{\min}}|,Y_{n_{\min}})$triangleright calculate starting MLE using (2)
5:$\mu_{n_{\min}}(x_{*})=\Sigma(x_{*},X_{n_{\min}}(x_{*}))\Sigma_{n_{\min}}^{-1}Y_{n_{\min}}$triangleright make prediction for $x_{*}$ using (4)
6:for $n\in\{n_{\min}+1,\dots,n_{\max}\}$ do
7:     $x_{n}=X_{n}(x_{*})\setminus X_{n-1}(x_{*})$triangleright newest point in neighborhood
8:     $\{\Sigma_{n}^{-1},\log|\Sigma_{n}|\}=\mathrm{update}(\Sigma_{n-1}^{-1},\log|\Sigma_{n-1}|,x_{n})$triangleright calculate new $\Sigma_{n}^{-1}$ and $|\Sigma_{n}|$ using (6)
9:     if $\mathrm{log}_{2}(n)\mod 1=0$ then
10:         $\{\tau^{2},\theta\}=\mathrm{MLE}(\Sigma_{n}^{-1},\log|\Sigma_{n}|,Y_{n})$triangleright recalculate MLE using (2)
11:     end if
12:     $\mu_{n}(x_{*})=\Sigma(x_{*},X_{n}(x_{*}))\Sigma_{n}^{-1}Y_{n}$triangleright make prediction for $x_{*}$ using (4)
13:end for
14:$\hat{n}(x_{*})=\underset{n\in\{n_{\min},\dots,n_{\max}\}}{\mathrm{argmin}}\mathrm{SE}\left(y_{*},\mu_{n}(x_{*})\right)$triangleright choose best neighborhood size using (5)
15:return $\mathrm{LAGP}_{\hat{n}(x_{*})}(x;X_{\hat{n}}(x),Y_{\hat{n}}))$triangleright make prediction for $x$ with $\hat{n}(x_{*})$
The details, summarizing the entirety of our approach to solving Eq. (5), is provided by Algorithm 1. Notice that the procedure ends after re-fitting LAGP with the chosen $\hat{n}(x_{*})$ with $x$ and all of $X_{N}$.
Continuing our illustration in Figure 4, the center panel provides a view of the SE-values encountered via Eq. (5) using the setup on the left, for $x=(-0.32,0.17)$, as $n$ ranges from $n_{\min}=12$ to $n_{\max}=160$. The result is $\hat{n}(x)=43$, and the indices for these locations are indicated with white text. Notice how $\hat{n}(x)$ is chosen such that $X_{\hat{n}}(x)$ is as big as possible while residing entirely in the red-response area. This process must be repeated for $x\in\mathcal{X}$ of interest, whereas only one is illustrated in the figure. Some of the details of how we recommend doing this are deferred to Section 5. A for loop around Algorithm 1 is one simple option, and is the basis of the illustration described next.
3.2 An Illustration
Now consider applying Alg. 1 independently on each $x\in\mathcal{X}$, e.g., testing on a $101\times 101$ grid, extending our running “Phantom” example. The right panel of Figure 4 uses color to indicate the predictive means obtained at each $x$ on the last line of the algorithm. Compared to the truth [Figure 3] the fit looks good, visually speaking, but not perfect. We show in Appendix B.1 that this fit is more accurate, in terms of RMSE, than an ordinary GP fit. Here we aim to illustrate the optimal choice of neighborhood size, $\hat{n}(x)$, in two ways. We begin with focus on two particular predictive locations: $(-0.32,0.17)$ and $(0.39,0.34)$, far from and near to the manifold of discontinuity, respectively. Observe how the optimal neighborhood size (OLAGP) adjusts as appropriate ($n=43$ far from the manifold; $n=14$ near).
Size OLAGP Modular JGP
12–20 4548 3914
21–40 2838 2853
41–60 745 933
61–80 414 697
81–100 319 298
101–120 139 396
121–140 511 544
141–160 687 566
Figure 5: Distribution of OLAGP and Modular JGP [Section 4] estimated neighborhood sizes.
Now consider how all 10K-odd neighborhoods, over the whole grid $\mathcal{X}$, compare to the laGP software default of $n=50$. Figure 5 shows how that choice (between the vertical lines) is not advantageous for the Phantom data. Many smaller, but also many much larger larger neighborhoods are estimated by OLAGP, and with much greater frequency. The figure also provides neighborhoods for the new method (“Modular JGP”) discussed next in Section 4.
Before turning to that, we note anecdotally that OLAGP performs worst at or near the manifold of discontinuity (See Appendix A.1). This is because it cannot choose $\hat{n}(x)$ small enough to exclude points from the opposite side of the manifold without maintaining a viable minimum ($n_{\min}=12$) for a GP fit. Observe how the neighborhood for $(0.39,0.34)$ contains several points in the red region. At $\hat{n}(x)=14$ for that location, the neighborhood can’t shrink much further. Consequently, the resulting mean predictions (4) are weighted averages of red- and yellow-region response values. This will yield predictions between 0.4 and 0.6, which we know from Figure 3 are inaccurate. Shrinking neighborhoods farther, even below $n_{\min}=12$, is a double-edged sword: less data means greater reactivity, but also lower information content. It would be better to allow a more flexible notion of neighborhood, beyond simple Euclidean distance in $X_{N}$. Next, we suggest how that can be done without fundamentally altering the LAGP setup by introducing a latent variable. To foreshadow, observe that the “Modular JGP” neighborhood sizes summarized in Figure 5 show fewer small neighborhoods, and fewer large neighborhoods compared to OLAGP.
4 Modular Jump Gaussian Processes
Next we target disparate levels of data from a jump process: first identifying marginally, then predicting spatially, ultimately constructing a feature that can be used as a predictor in a (LA)GP.
4.1 Identifying and predicting levels
Data $Y_{N}$ from a jump process can often be partitioned marginally (i.e., $y$ without $x$), which can inform a more complex spatial partition over $x$. To see this, consider again the histogram of the marginal $Y_{N}$ for the Phantom data [Figure 3, right]. There are two, roughly symmetric, well-separated densities: centered at 0.2 and 0.8. These modes are so well separated you can eyeball it, say at $y=0.5$.
Visual partitioning of the marginal response will not always be possible, as we shall show momentarily, and it is unsatisfying to require a human in the loop of an otherwise easily automatized procedure. We assume generally that data from jump processes can be captured, marginally, as a mixture distribution with two components. Under this assumption, each observation $y_{i}$ is associated with a latent variable $\ell_{i}\in\{1,2\}$ describing the component $y_{i}$ belongs to. Of interest to us is the quantity $\mathbb{P}(\ell_{i}=j)$ for $j=1,2$, the probability that observation $y_{i}$ belongs to component $j$.
A model-based clustering method can be used to identify the level labels of each observation in $Y_{N}$. We prefer EM (Dempster et al.,, 1977; McLachlan and Krishnan,, 1997) under Gaussian mixtures (McLachlan and Basford,, 1988, Chapter 2), extracting $\mathbb{P}(\ell_{i}=j)\equiv\tau_{ji}$ in the notation of that paper. Then, let $c_{i}=\underset{j\in\{1,2\}}{\arg\max}\,\mathbb{P}(\ell_{i}=j)$, the highest probability component for $y_{i}$. One view of the results of such a clustering is indicated via the marginal histogram in Figure 3, where color indicates each $c_{i}$, $i=1,\dots,n$; bars colored red are sorted into component 1, yellow into component 2.
Figure 6: Left: Star histogram; Middle: Modified Michalewicz function histogram; Right: AMHV transport histogram. Colors represent the two levels identified by each clustering algorithm.
Figure 6 provides similar views for three other examples entertained in our empirical work in Section 5, including our motivating real-data set (right panel). In these plots, we also include lines showing the two-component Gaussian mixture estimated by EM. Like the Phantom, the Star example (left panel) is relatively simple. Visual partitioning would have sufficed. The other two benefit from a deliberate, model-based approach. In both of these examples (middle and right panels), it may well be that more than two components (or non-Gaussian ones) would be preferable. However, we aim to show that a crude marginal “partition” is powerful when this information is reincorporated spatially, which we describe next.
Let $c_{i}\in\{1,2\}$, $i=1,\dots,N$ denote the level learned for each $y_{i}\in Y_{N}$ via model based clustering, with vector $C_{N}$ holding all $N$ levels. For example, we could have $c_{i}=2$ correspond to “yellow” and 1 for “red”, or vice-versa. Then, pair these with inputs $X_{N}$ to form $(X_{N},C_{N})$, comprising “data” for a spatial classification task. Here we shall write $\hat{f}(\cdot)=\mathrm{class}(X_{N},C_{N})$, generically, for the predictive equations of a fitted, spatial classification. One option is (basis-expanded) linear logistic regression, or logistic GP classification (e.g Rasmussen and Williams,, 2006, Chapter 3). Any accurate method will do; our preferred implementation is detailed in Section 5. We may then use $\hat{f}$ to predict component membership at testing locations $\mathcal{X}$: $\hat{f}(x)=\mathbb{P}(y(x)=1)$ for any $x\in\mathcal{X}$. Predicted probabilities for the Phantom data are shown in the left panel of Figure 7.
Figure 7: Left: Jump feature. Center: Nearest 160 neighbors to $x=(-0.32,0.17)$ after adding the jump feature. White points are closest to $x$, and black points are furthest away. Right: Modular JGP fit. Blue dots are two predictive locations and the surrounding black dots are the selected neighborhood.
The other panels will be discussed momentarily, in Section 4.2. Observe how the contours in this plot are similar in shape to the original data [see Figure 3], with values close to 0 in deeply yellow regions and values close to 1 in deeply red regions. In particular, note that $\hat{f}(x)=0.5$ for locations $x$ near the manifold of discontinuity, where component membership is least certain.
4.2 A new feature for GP regression
Now recall that stationary kernels enforce a covariance structure that is based only on the pairwise distances between points (3). As the distance between two points increases, their covariance decreases. This covariance is then used in the GP mean prediction (4), where it serves as a weight. This means that points that are close to the predictive location have a lot of influence on its mean, and points that are far away have little influence. Therefore, if we can find a way to modify the distances between points so that points from the same component are closer together than those from opposite components, this covariance structure would naturally place more influence on points from the same component than those from the opposite one. A simple way to modify these distances without sacrificing the original distance structure is by augmenting with a new input feature (Bornn et al.,, 2012).
This new variable should be related to component membership. We propose using the probability of component membership [Section 4.1], which should be high for observations likely to be in one component and low for the other. E.g., classifier $\hat{f}(x)$ predicts the probability each location $x$ belongs to the red region. I.e., it is large (close to 1) in deeply red regions and small (close to 0) in deeply yellow regions. Suppose we treat those values as new feature, comprising a new column of $X_{N}$ for GP modeling. In the new, higher dimensional input space, points in the red region will be closer together than those from the yellow region, and vice versa. Under the likelihood (2) and when predicting (4), points from the same component will have more influence on each other than on points from opposite components.
For this to work, we need to know the probability of component membership for both $Y_{N}$ and $\mathcal{Y}$. Section 4.1 provides estimates of $C_{N}$, for $Y_{N}$, and $\hat{f}(\mathcal{X})$, for $\mathcal{Y}$. The issue is that $C_{N}$ is not a probability; it is a vector consisting of 1s and 2s describing the most likely component a point in $Y_{N}$ belongs to. However, $\hat{f}(\cdot)$ may be used to predict, in-sample, the probability that $Y_{N}$ belongs to a particular component. Vectors $\hat{f}(X_{N})$ and $\hat{f}(\mathcal{X})$ comprise the training and testing jump features for $X_{N}$ and $\mathcal{X}$, respectively. These may be appended as new columns $X\equiv[X,\hat{f}(X)]$ and $\mathcal{X}\equiv[\mathcal{X},\hat{f}(\mathcal{X})]$, overloading the notation somewhat.
To illustrate the effect this new feature has, consider the center plot of Figure 7, again focusing on $x=(-0.32,0.17)$ and its nearest neighbor $x_{*}$. We also plot the 160 NNs to $x$, color coded by the rank of their distance from $x$. The points that are the lightest are closest to $x$, and those that are the darkest are furthest away. Note that these distances are calculated prior to hyperparameter fitting. Observe first that the neighborhood is no longer circular; it has become oval-shaped to avoid the yellow region. Any numbered points that are in the yellow region are much further away than their 2d distance would suggest. Their colors are darker than nearby points the red region, and thus farther from $x$.
We generically notate the fitting stage as $\mathrm{GP}([\mathcal{X},\hat{f}(\mathcal{X})];[X_{N},\hat{f}(X_{N})],Y_{N})$. where “GP” can stand in for any type (e.g., global, LAGP, OLAGP) depending on the problem at hand. We refer to this method, beginning with the clustering algorithm and ending with the final GP fit, as a “Modular JGP”. Algorithm 2 provides the details for quick reference.
Algorithm 2 Outline of Modular JGP
1:$X_{N},Y_{N},\mathcal{X}$
2:$\hat{C}_{N}=\text{clust}(Y_{N})$triangleright perform clustering algorithm on marginal $Y_{N}$, yielding $C_{N}\in\{1,2\}^{N}$
3:$\hat{f}(\cdot)=\text{class}(X_{N},C_{N})$triangleright build classifier mapping $X_{N}$ to $C_{N}$, yielding $\hat{f}(x)=\text{Pr}(y(x)=1)$
4:$\text{GP}([\mathcal{X},\hat{f}(\mathcal{X})];[X_{N},\hat{f}(X_{N})],Y_{N})$triangleright fit any GP (e.g. OLAGP) with jump feature
The right panel of Figure 7 provides a Modular JGP fit on the Phantom data, contrasting with Figure 4. In this case the GP used is an OLAGP, chosen to illustrate the change in neighborhood structure induced by the Modular JGP. Here, in Figure 7, both neighborhoods are smaller, but are completely contained to the same side of the manifold. The Modular JGP has visually improved prediction both visually and via the root mean square error (RMSE; see Appendix B.1) over the global GP, LAGP, and OLAGP.
Now return to Figure 5, summarizing all 10K-odd $\hat{n}(x)$ for OLAGP and Modular JGP. Again, the choice of $n=50$ (the LAGP default) is not particularly advantageous, except for a small percentage of $x\in\mathcal{X}$. Notice that the Modular JGP slightly prefers more mid-range neighborhood sizes than the optimal neighbors GP. In particular, many fewer small $\hat{n}(x)$ are chosen since the neighborhood shape is no longer circular in the original input space (i.e., ignoring the new feature). This facilitates excluding points from the opposite component without resorting to extremely small neighborhood sizes.
Our Modular JGP is purposefully generic: clustering, classifying, predicting, and then fitting with a new feature. We mentioned some of the possible choices for each step (e.g. EM, CGP, and OLAGP), but any methods for clustering, classifying, and regression are fair game. The best choices are likely context/data-dependent. We aim to show that our preferred instance of Modular JGP is an improvement on existing methods in the context of surrogate modeling of jump processes from the recent literature.
5 Implementation and empirical benchmarking
Here we present three simulated and one real-world example(s) to demonstrate the effectiveness of OLAGP and the Modular JGP. A summary of each data set/experiment is included in Table 1. Details are provided in separate sub-sections to follow. Broadly, for each data set, we conduct a Monte Carlo (MC) experiment which repeatedly randomly partitions the data into 90:10 training and testing sets. In each each repetition, we sample a random subset of 90% of the data set to be used as the training set, and use the remaining 10% as the testing set. The number of MC trials depends on the computational expense implied by the training data set size $n$, which varies in the experiments. We present five competing methods:
  • Global GP via newGPsep in laGP (Gramacy,, 2016), using defaults. This was excluded from the Modified Michalewicz experiment due to size constraints.
  • LAGP via aGPsep in laGP (Gramacy,, 2016), with method = "nn" and all other defaults.
  • OLAGP from Section 3 with $n_{\min}=12$ and $n_{\max}=160$.
  • Original JGP via Park, (2022)’s MATLAB library using classification EM and quadratic boundaries.
  • Modular JGP from Section 4 with further details provided in Table 1.
All experiments were completed in R. Code is provided at https://bitbucket.org/gramacylab/jumpgp. The Original JGP leveraged R.matlab (Bengtsson,, 2022).
Data $N$ $p$ Clustering Classifier GP Type Reps
Phantom 10,201 2 EM CGP OLAGP 100
Star 10,201 2 EM CGP OLAGP 100
Modified Michalewicz 100,000 4 EM Random Forest OLAGP 30
AMHV transport 9,503 4 EM CGP Global GP 100
Table 1: Implementation summary for each data set/experiment.
OLAGP is parallelized over predictive locations $x\in\mathcal{X}$, turning statistical independence into computational exclusivity via foreach (Microsoft and Weston,, 2022). For the Modular JGP, we utilize EM-based clustering and either a CGP (Williams and Barber,, 1998) or a random forest (Breiman,, 2001) for classification via mclust (Scrucca et al.,, 2023), kernlab (Karatzoglou et al.,, 2004) and randomForest (Liaw and Wiener,, 2002), respectively. We only prefer a random forest over a CGP when $n$ is large, for example during the Modified Michalewicz experiment. Again, see Table 1. We measure model performance out-of-sample using $\mathrm{RMSE}=(\frac{1}{m}\sum_{x\in\mathcal{X}}(y(x)-\mu(x))^{2})^{1/2},$ for which lower is better.
5.1 Phantom
In Sections 34, we assessed model performance (visually and with the RMSE) on the Phantom data for a single train–test partition. The outcome of a full MC experiment is summarized by boxplots in the left panel of Figure 8.
Figure 8: Left: MC experiment on the Phantom data set; Right: A line plot showing the RMSE at individual runs for each of the five methods. The Modular JGP performs best in 96 of 100 runs.
For a more granular view, the right panel draws black lines conneting RMSEs from the same rep, allowing us to see the best performing method for every training/testing set combination. We conclude that most of the overlap of the boxplots is due to variability in the training/testing split. The ordering of accuracy from one MC rep to the next is consistent across methods.
In these plots (and all plots in this section), our two methods are shown third from right (OLAGP) and furthest right (Modular JGP). Both show improvement over the global GP and LAGP, although OLAGP is typically outperformed by the Original JGP. This is unsurprising, since the Original JGP was designed specifically for data from jump processes, and OLAGP was not. But notice that the Modular JGP consistently outperforms the Original JGP, for this data set and broadly over the next few subsections. Looking first at the line plot on the right of Figure 8, we see that for almost every data set (in particular, 96 of 100), predictions made by the Modular JGP are more accurate than those made by the Original JGP. Additionally, the boxplot shows that the Original JGP has higher variability than the Modular JGP.
5.2 Star
Now consider the “Star” data, which we created for this paper. The goal was to have something like the Phantom, but more angular. Details for how the data set was created are provided in Appendix A.2. See the left panel of Figure 9, with marginal response histogram depicted earlier in the left panel of Figure 6. Like the Phantom, the Star has two well-separated regimes that form a clear shape on the response surface. But the Star is more difficult to model because all points in the yellow regime are geographically close to orange ones. It is hard to build a neighborhood inside the yellow portion of the star.
Figure 9: Model fit and chosen neighborhoods for Star data set. Left: Ground truth; Center: OLAGP fit of Star data set; Right: Modular JGP fit of Star data set. Two predictive locations, in blue, are shown with their chosen neighborhoods, in black.
Repeating themes from the Phantom analysis, consider model performance for a single Star train–test split. Mean fits for the global GP and LAGP are shown in Figure 16 and in Appendix B.2, whereas those of OLAGP and the Modular JGP are shown here in the right two panels of Figure 9. RMSEs can be found in Appendix B.2, where we see that the local methods perform well, with the Modular JGP producing the best predictions of the four. The scenarios depicted in right two panels focus on $x=(0.15,0.8)$ and $(0.83,0.11)$. Compared to LAGP, OLAGP has chosen a neighborhood of roughly the same size for $(0.83,0.11)$, and much smaller for $(0.15,0.8)$; however, this smaller neighborhood still contains points from the orange regime. The Modular JGP provides an interesting change: a much smaller neighborhood size for $(0.15,0.11)$, and a larger neighborhood size for $(0.83,0.11)$. Both neighborhoods have adapted their shape to only include points from the same regime, and this improves prediction overall.
Figure 10: Left: MC experiment on the Star data set; Right: A line plot showing the RMSE at individual runs for each of the five methods. The Modular JGP performs best in 76 of 100 runs.
Figure 10 summarizes a MC experiment along similar lines. Here again, OLAGP outperforms both the global GP and LAGP; the two JGP methods typically outperform OLAGP. The Original JGP is again highly variable. Although it outperforms the Modular JGP more frequently than for the Phantom (Original JGP was better in 24 of 100 runs), the Modular JGP is still the best performing method for the Star.
5.3 Modified Michalewicz
For a higher-dimensional example, we adapted the so-called Michalewicz function, which can be found in the Virtual Library of Simulation Experiments (Surjanovic and Bingham,, 2013). Details can be found in Appendix A.3. Michalewicz is dimension-adaptive; here we present a 4d analysis to match the real data analysis coming in Section 5.4. An LHS design is used with $N=100,000$ training inputs. The marginal distribution of the response is shown in the center panel of Figure 6. Unlike the previous two data sets, the response is not plausibly composed of two Gaussian components. Instead, it is comprised of a point mass around 0.5 with other points scattered between in $[-4,0]$. Nevertheless, we find that EM with two Gaussian components works well in the Modular JGP setup. One goal of the presentation here is to demonstrate how even a naïve marginal clustering often works well in practice.
Figure 11: Left: MC experiment on the Modified Michalewicz data set; Right: A line plot showing the RMSE at individual runs for each of the four methods. The Modular JGP performs best in 27 of 30 runs.
Since $N$ here is very large, we must omit the global GP as a compactor in this experiment. Instead of a CGP we deploy a random forest as our classifier instead. Finally, we limit the number of MC reps to thirty. Results are provided in Figure 11. LAGP, optimal neighbors JGP, and the Modular JGP all considerably outperform the Original JGP. LAGP and OLAGP perform comparably, although interestingly LAGP seems to slightly outperform OLAGP in this case. The Modular JGP performs best of all, yielding the lowest RMSE in 27 of the 30 reps.
5.4 AMHV transport
Finally, consider our motivating AMHV transport data. Specifically, we used the larger version of the data set described in Section 5.2 of Park et al., (2023). The input dimension is $d=4$ and $N=9{,}500$. Although $N$ in moderately large in absolute terms, 4d is a large volume to fill. So, in a certain sense, this example is the most sparse of the ones entertained in this paper. Outputs are derived from a simulation of the behavior of AMHVs in a particular facility, tallying the average transport time of the AMHVs, including any time spent waiting for the vehicle to arrive. Park et al., (2023) show that as production in the facility increases, transport time, which depends on various design variables, increases and becomes a major bottleneck to the entire manufacturing process. The four varied here include vehicle speed, vehicle acceleration, minimum distance required between vehicles, and the maximum search range of empty vehicles.
A view of the marginal distribution of the response for these data was provided earlier, in the right panel of Figure 6. This is the first example we have looked at where the two clusters overlap. Additionally, the sparse nature of this data set means that local methods do not perform nearly as well. (The closest points in $X_{N}$ to a predictive location $x$ may actually be far and not representative.) This is a particular problem for OLAGP, which uses $x_{*}$ as a proxy for $x$. Consequently, we prefer a global rather than local GP in our application of the Modular JGP for these data.
Figure 12: Left: MC experiment on the AMHV transport data set; Right: A line plot showing the RMSE at individual runs for each of the five methods. The Modular JGP performs best in 95 of 100 runs.
The results of the experiment are shown in Figure 12. As expected, the local methods do not perform as well as the global ones, with OLAGP suffering most. But the global GP performs well, producing comparable results to the Original JGP. The Modular JGP produces only slightly better results, but is consistent in doing so (lowest RMSEs in 95 of the 100 reps).
6 Discussion
We introduced two ideas, adapting LAGP neighborhoods locally (OLAGP) learning a spatial feature from marginal clustering. Separately, both improve predictions for data from jump processes. They shine brightest when used together. We call this hybrid Modular JGP, since it involves many of the same ingredients as the Original JGP, except with off-the-shelf subroutines daisy-chained together.
One opportunity for future work pertains specifically to OLAGP. We mentioned previously that this method has thus far only been implemented for neighborhoods chosen using NN. However, it has been shown that NNs can be suboptimal compared to other subdesigns in the GP contexts (Vecchia,, 1988; Stein et al.,, 2004). Gramacy and Apley, (2015) demonstrated that neighborhoods chosen using via design criteria mean-square prediction error (MSPE) often yield superior LAGP predictions. We expect that to extend to the OLAGP context, but investigating that represents future work.
Other potential opportunities target the Modular JGP more holistically. The first involves extending the Modular JGP to noisy functions. The main concern in this case is whether the signal-to-noise ratio is large enough for the GP to accurately capture jumps in the data. The second involves extending the Modular JGP to allow more than two jump levels. Here we assumed that the marginal response of data from jump processes could be described as a mixture distribution with two components. However, it is possible for the response to have more than two levels. In that case, it might help to augment with a new feature for each added level. In other words, for $\ell$ levels we require $\ell-1$ new features. This would require a multiclass classification subroutine, however both of the methods we used – classification GP from kernlab and random forest from randomForest – have this capability.
Acknowledgements
ARF and RBG are grateful for support from NSF 2152679. CP is grateful for support from NSF 2420358. Part of this research was initiated within the frame of the consortium in Applied Mathematics CIROQUO, gathering partners in technological research and academia in the development of advanced methods for Computer Experiments.
Declaration of Interest
The authors report no conflict of interest.
SUPPLEMENTARY MATERIAL
Appendix A Data Generation Details
Here we describe the process for generating the Phantom, Star, and Modified Michalewicz data sets.
A.1 Phantom
The Phantom data set is generated in the same way as the Phantom data set in Park, (2022), but with different hyperparameter values. This version of the Phantom is evaluated on an evenly spaced $101\times 101$ grid in $[-0.5,0.5]^{2}$. The response is generated using the partitioning scheme shown in the left panel of Figure 13. In this plot, the color represents the true component membership of every input combination: white pixels belong to component 1, and black pixels belong to component 2. Data are drawn from a $\mathcal{N}_{N_{i}}(\mu_{i},\Sigma(X_{N_{i}}))$, where $N_{i}$ is the number of observations in component $i$, $\mu_{i}$ is the mean of component $i$, and $X_{N_{i}}$ is the subset of observations which belong to component $i$. The mean vector is $\mu=[27,0]^{\top}$ and the covariance matrix $\Sigma(X_{N_{i}})$ is calculated using the exponential covariance function (3) with $\tau^{2}=9$ and $\theta=[0.1,0.1]^{\top}$. After sampling all the data, the response was standardized so that $Y_{N}\in[0,1]$.
Figure 13: Left: Image describing component membership for Phantom data set; Right: Image describing component membership for Star data set.
A.2 Star
The Star data set is inspired by the Star data set in Park, (2022), but uses a different image for its partitioning structure. This version of the Star is evaluated on an evenly spaced $101\times 101$ grid in $[0,1]^{2}$. The response is generated using the partitioning scheme shown in the right panel of Figure 13. Like with the Phantom, the color represents the true component membership: white for component 1, and black for component 2. Data are drawn from a $\mathcal{N}_{N_{i}}(\mu_{i},\Sigma(X_{N_{i}}))$. The mean vector is $\mu=[-10,10]^{\top}$ and the covariance matrix $\Sigma(X_{N_{i}})$ is calculated using the exponential covariance function (3) with $\tau^{2}=3$ and $\theta=[0.01,0.01]^{\top}$. After sampling all the data, the response was standardized so that $Y_{N}\in[0,1]$.
A.3 Modified Michalewicz
The Modified Michalewicz is a modification of the Michalewicz function from the Virtual Library of Simulation Experiments (Surjanovic and Bingham,, 2013). The original Michalewicz function is defined as $f(x)=-\sum_{i=1}^{d}\sin(x_{i})\sin^{2m}\bigl{(}\frac{ix_{i}^{2}}{\pi}\bigr{)}$ for $x\in[0,\pi]^{4}$ and $d$ in arbitrary dimension. We use $d=4$ and $m=10$ (the recommended default) with inputs coded into $[0,1]^{4}$. We first evaluate $f(x)$ using an LHS of size 100,000. The Modified Michalewicz is then defined as:
$y(x)=\begin{cases}f(x)&f(x)\leq-0.01\\ f(x)+0.5&f(x)>-0.01.\\ \end{cases}$ (7)
We use this modification because the response surface of $f(x)$, without modifications, contains no jumps. It is comprised of many local minima that are surrounded by a nearly flat surface around 0. Consequently, the marginal response is comprised of two “levels”: a point mass around 0, and the rest of the surface. To create the jumps (i.e. discontinuities) that we expect, we raise the flat area of the surface by 0.5. This creates a manifold of discontinuity where the response jumps from the point mass to the rest of the surface.
Appendix B Global and LAGP Fits
We provide the mean fits given by a global GP and LAGP for the Phantom and Star data sets to highlight the improvements made by OLAGP and the Modular JGP. We also provide the difference between the predicted mean fit and the truth for each data set.
B.1 Phantom
The left two panels of Figure 14 show the predicted mean fits for the Phantom data set using global GP and LAGP models.
Method RMSE
Global GP 0.1543
LAGP 0.1514
OLAGP 0.1316
Modular JGP 0.1207
Figure 14: Left: Global GP fit. Center: LAGP fit. The blue dots are two predictive locations and the surrounding black dots are the selected neighborhood for that predictive location. Right: RMSE of Global GP, LAGP, OLAGP, and Modular JGP fits.
Both fits are visually less accurate than the OLAGP fit in Figure 4 and Modular JGP fit in Figure 7. This is verified by the table of RMSE values in the right panel, which shows that both OLAGP and the Modular JGP produce more accurate predictions than either method. The LAGP fit in the center panel also shows the selected neighborhoods for the same two predictive locations used in Sections 3 and 4, $(-0.32,0.17)$ and $(0.39,0.34)$. Notice in particular that the neighborhood for $(0.39,0.34)$ contains points in both the orange and yellow regions, which explains why our prediction is inaccurate. We are able to select smaller and differently shaped neighborhoods using OLAGP and Modular JGP so that our predictions are more accurate. We can verify that OLAGP and the Modular JGP make more accurate predictions by looking at Figure 15, which shows the absolute difference between the mean predictions and the truth.
Figure 15: Left: Difference between Global GP fit and truth; Center Left: Difference between LAGP fit and truth; Center Right: Difference between OLAGP fit and truth; Right: Difference between Modular JGP fit and truth.
Notice that the difference is largest along the manifold of discontinuity.
B.2 Star
The left two panels of Figure 16 show the predicted mean fits for the Star data set using global GP and LAGP models.
Method RMSE
Global GP 0.2716
LAGP 0.1967
OLAGP 0.1599
Modular JGP 0.1462
Figure 16: Left: Global GP fit. Center: LAGP fit. The blue dots are two predictive locations and the surrounding black dots are the selected neighborhood for that predictive location. Right: RMSE of Global GP, LAGP, OLAGP, and Modular JGP fits.
Both fits are visually less accurate than the OLAGP and Modular JGP fits in Figure 9. The global GP performs particularly bad in this case, likely because the proportion of observations in component 1 is much larger than the proportion of values in component 2. The table of RMSE values in the right panel verifies our intuition, showing that OLAGP and the Modular JGP produce more accurate predictions than either method, particularly the global GP. The LAGP fit in the center panel also shows the selected neighborhoods for the same two predictive locations used in Section 5.2, $(0.15,0.8)$ and $(0.83,0.11)$. Notice in particular that the neighborhood for $(0.15,0.8)$ contains points in both the orange and yellow regions, which is why our prediction is inaccurate. We are able to select smaller and differently shaped neighborhoods using OLAGP and Modular JGP so that our predictions are more accurate. We can verify that OLAGP and the Modular JGP make more accurate predictions by looking at Figure 17, which shows the absolute difference between the mean predictions and the truth.
Figure 17: Left: Difference between Global GP fit and truth; Center Left: Difference between LAGP fit and truth; Center Right: Difference between OLAGP fit and truth; Right: Difference between Modular JGP fit and truth.
Notice that the difference is largest along the manifold of discontinuity.